# xMIL: Insightful Explanations for Multiple Instance Learning in Histopathology

### v1

@@ A.3 Histopathology experiments: data and training details
A.4 Faithfulness evaluation: patch flipping
We binned the patches of slide $X$ into 100 ordered groups $(\mathbf{r}_{1},\cdots,\mathbf{r}_{100})$, where $\mathbf{r}_{i}$ is the set of all patches whose attribution scores are between the $(100-i-1)$-th and $(K-i)$-th percentiles of $\{\mathcal{H}(X,F,f_{1}(\mathbf{x}_{k}))\}_{k}$, for example, $\mathbf{r}_{1}$ is the set of the most relevant 1% patches of $X$.
Patch dropping. Following the region perturbation strategy introduced in [67], we progressively excluded the most relevant regions from slide $X$, i.e. in the $k$-th iteration, the most relevant $k$of patches were excluded. Formally, the perturbation procedure can be formulated as the following:
$\begin{gathered}X_{\text{morf}}^{(0)}=X\\ X_{\text{morf}}^{(k)}=\mathcal{P}(X,k)\\ \mathcal{P}(X,k)=\bigcup_{i=k+1}^{100}\mathbf{r}_{i}\end{gathered}$ (3)
where $\mathcal{P}(X,k)$ is a perturbation function that excludes the most relevant $k$of patches from slide $X$. Note that at step $k=100$ all the patches are excluded and therefore, $X_{\text{morf}}^{(100)}=\mathbf{0}$, where $\mathbf{0}$ is an array of zeros.
Comparing heatmaps. We define the quantity of interest for the comparison of different heatmaps as the area under the perturbation curve (AUPC) as follows:
$\text{AUPC}(X,\mathcal{H})=\frac{1}{100}\sum_{k=0}^{100}f(X_{{\text{morf}}}^{(k)})$ (4)
@@ A.4 Faithfulness evaluation: patch flipping
We ran all the AttnMIL experiments either on an A100 40GB GPU and the TransMIL experiments on a single CPU.
A.5 Extracting insights from xMIL-LRP heatmaps: additional results
@@
We display further exemplary heatmaps of TransMIL model predictions in the HNSC HPV dataset in Figures 6 and 7.

### v2

@@ A.4 Histopathology experiments: Data and training details
A.5 Faithfulness evaluation: Patch flipping
Given a slide $X=\{\mathbf{x}_{k}\}_{k=1}^{K}$ and a heatmaping function $\mathcal{H}$ producing the explanation scores of the instances in $X$, i.e., $\mathcal{H}(\mathbf{x}_{k})=\hat{\epsilon}_{k}$, we binned the patches of slide $X$ into 100 ordered groups $(E_{1},\cdots,E_{100})$, where $E_{i}$ is the set of all patches whose attribution scores are between the $(100-i)$-th and $(100-i+1)$-th percentiles of the explanation scores of the instances of $X$, for example, $E_{1}$ is the set of the most relevant 1% patches of $X$ and $E_{100}$ is the least relevant 1% of the patches.
Patch dropping. Following the region perturbation strategy introduced in [74], we progressively excluded the most relevant regions from slide $X$, i.e. in the $n$-th iteration, the most relevant $n$of patches were excluded. Formally, the perturbation procedure can be formulated as the following:
$\begin{gathered}X_{\text{morf}}^{(0)}=X\\ X_{\text{morf}}^{(n)}=\mathcal{P}(X,n)\\ \mathcal{P}(X,n)=\bigcup_{i=n+1}^{100}E_{i}\end{gathered}$ (11)
where $\mathcal{P}(X,n)$ is a perturbation function that excludes the most relevant $n$of patches from slide $X$. Note that at step $n=100$ all the patches are excluded and therefore, $X_{\text{morf}}^{(100)}=\emptyset$, where $\emptyset$ is an empty set, for which we pass an array of zeros to the model.
Comparing heatmaps. We define the quantity of interest for the comparison of different heatmaps as the area under the perturbation curve (AUPC) as the following:
$\text{AUPC}(X,\mathcal{H})=\frac{1}{100}\sum_{n=0}^{101}f(X_{{\text{morf}}}^{(n)})$ (12)
where $f$ is the model’s function.
@@ A.5 Faithfulness evaluation: Patch flipping
We ran all the AttnMIL experiments on an A100 40GB GPU and the TransMIL experiments on a single CPU.
A.6 Faithfulness evaluation: Additional results
We present the patch dropping results for AttnMIL in Figure 4.
Figure 4: Patch dropping results for AttnMIL. The first row depicts the perturbation curves, where the solid lines are the average perturbation curve and the shaded area is the standard error of the mean at each perturbation step. Each boxplot on the second row shows the distribution of AUPC values for all test set slides per explanation methods. In each boxplot, the red line marks the median and the red dot marks the mean. Lower perturbation curves and AUPCs represent higher faithfulness.
A.7 Extracting insights from xMIL-LRP heatmaps: Additional results
@@
We display further exemplary heatmaps of TransMIL model predictions in the HNSC HPV dataset in Figures 6, 7, and 8.